# Better Safe Than Sorry? Overreaction Problem of Vision Language Models in Visual Emergency Recognition

### v2

@@ Abstract
Vision-Language Models (VLMs) have shown capabilities in interpreting visual content, but their reliability in safety-critical everyday life scenarios remains insufficiently explored. We introduce VERI (Visual Emergency Recognition Dataset), a diagnostic benchmark comprising 200 images organized into 100 contrastive pairs. Each emergency scene is paired with a visually similar but safe counterpart through human verification and refinement. Using a two-stage evaluation protocol—risk identification and emergency response—we assess 14 VLMs (2B to 124B parameters) across medical emergencies, accidents, and natural disasters. Our analysis reveals an ”overreaction problem,” where models accurately identify genuine emergencies (70-100% success rate) but produce high false-positive rates, misclassifying 31-96% of safe situations as dangerous. Ten safe scenarios were universally misclassified by all models regardless of scale. This ”better-safe-than-sorry” bias primarily results from contextual overinterpretation (88-93% of errors), challenging VLM reliability in safety-critical applications. These findings highlight fundamental limitations in current VLM architectures, which persist despite increased model scale. Our results demonstrate an urgent need for strategies specifically improving contextual reasoning in ambiguous visual situations. The consistently low performance of the model indicates that these data serve effectively as a diagnostic data set.
@@ 1 Introduction
Vision-Language Models (VLMs) have made remarkable progress in understanding visual content, advancing from simple object recognition to sophisticated scene understanding and contextual reasoning [17, 10, 4]. These models now power a wide range of applications from content moderation to assistive technologies [26]. However, their reliability in safety-critical scenarios remains underexplored. This gap raises a critical question with real-world implications: To what extent can current VLMs distinguish between genuine emergencies and visually similar but safe scenarios? To investigate this question systematically, we utilized generated synthetic (fake) data, which allows us to create controlled comparisons between dangerous and benign situations while maintaining visual similarity.
False visual perceptions of emergencies can lead to costly resource mobilization. In October 2023, firefighters were dispatched to two separate incidents in New York and Seoul. In both cases, they discovered that high-definition fireplace videos playing on television screens had been mistaken for actual fires [15, 23]. Such false alarms carry substantial costs. Each incident costs approximately \$1,000-2,500, with annual costs exceeding \$1 billion in the United States alone [19]. As VLMs power applications from smart home monitoring to CCTV surveillance, understanding their tendency to ”overreact” becomes crucial [24]. As VLMs increasingly power applications from smart home monitoring to CCTV surveillance, understanding their tendency to ’overreact’ becomes crucial  [24]. Current research on safety-critical VLMs has primarily focused on specialized domains. These include medical diagnosis systems [6], autonomous driving perception [7], and industrial robotics safety mechanisms [8]. Our work addresses a distinct gap: everyday emergency recognition across domains. This task requires different contextual reasoning capabilities—distinguishing between visually similar scenarios based on subtle contextual cues rather than domain-specific features. Existing benchmarks typically evaluate models on isolated images, rarely using contrastive pairs that test the ability to distinguish between visually similar but semantically distinct situations.
To address these gaps, we introduce VERI (Visual Emergency Recognition Dataset), a diagnostic benchmark of 200 images comprising 100 contrastive pairs. Inspired by compact yet influential diagnostic benchmarks such as Winogender (720 examples) [18], WSC (60 examples) [13], and PAIRS (200 images) [9], VERI prioritizes diagnostic value over quantity. We employ controlled image pairs to evaluate VLMs’ ability to distinguish between emergency and non-emergency situations. Each pair underwent over 5 rounds of iterative refinement and verification to balance visual similarity with semantic distinction.
@@
  • We introduce VERI, a benchmark with a two-stage protocol assessing risk identification and action suggestion in emergency scenarios.
  • Through evaluation of 14 VLMs (2B-124B parameters), we reveal a systematic ”overreaction problem” persisting across architectures and scales, with 10 safe scenarios misclassified by all models.
  • We identify two error patterns (contextual and visual misinterpretations), with contextual errors dominating (88-93%) and varying by emergency category.
  • We show that the overreaction problem persists despite model scaling, requiring targeted architectural improvements for contextual emergency assessment.
@@ 2.1 Dataset Design and Taxonomy
Effective emergency detection requires distinguishing between genuine threats and visually similar but safe situations. To address this challenge, the VERI dataset (200 images) has been specifically designed to study risks in everyday life, rather than focusing on specialized domains such as medical diagnosis or industrial safety. This focus on everyday risks is particularly relevant given the widespread adoption of autonomous driving and AI-driven robotics in our daily lives. By establishing this diagnostic benchmark, our study deliberately narrows its research scope to typical daily-life scenarios, where accurate threat detection is becoming increasingly critical. The core design principle of VERI is the contrastive pair approach—each entry consists of two images with high visual similarity but fundamentally different semantic implications: one depicting a genuine emergency requiring intervention and the other showing a visually similar but safe scenario. As illustrated in Figure 2, these pairs include contrasts such as actual medical emergencies versus training simulations, genuine accidents versus staged scenarios, and real disasters versus controlled environments. This contrastive structure enables direct assessment of a model’s ability to distinguish between superficially similar scenes with different safety implications. We organized VERI into three major categories of emergency situations (Accidents & Unsafe Behaviors, Personal Medical Emergencies, and Natural Disasters), as shown in Table 1. Each category presents unique visual challenges, from subtle physiological distress cues to complex environmental interpretation, as demonstrated in Figure 2.
@@ 2.2 Image Creation Process
Creating visually similar but semantically distinct image pairs required careful design to balance realism with controlled visual properties. We employed a multi-stage approach:
@@
Stage 3: Refinement and Validation. We iteratively refined images using inpainting until they matched our requirements (averaging 5+ cycles per pair). Three evaluators (two researchers and an independent annotator) validated each image against our criteria. The process continued until unanimous agreement was reached on all criteria (Cohen’s $\kappa=1.00$, $n=200$ pairs, $600$ decisions), eliminating annotation noise. These criteria ensured emergency scenes clearly required intervention, safe scenes were unambiguously non-threatening, and pairs maintained appropriate visual similarity.
@@
Our process prioritized diagnostic value over quantity, focusing on expert-guided refinement to effectively challenge VLMs’ contextual reasoning. Because data for everyday emergency recognition did not previously exist, generating a large volume of realistic images by abstracting from existing raw data was impractical. Moreover, for diagnostic benchmarks with specific evaluation goals, the qualitative value and the control of variables are significantly more important than sheer quantity. This principle aligns with other influential diagnostic benchmarks in the literature. For instance, the PAIRS benchmark comprises 200 images, and the study ’Judging LLM-as-a-Judge with MT-Bench and Chatbot Arena’ (Zheng et al., 2023) evaluated models using only 80 carefully constructed questions across eight fields.
@@ 3.1 Models
We evaluated 14 open-source VLMs (2B-124B parameters) across different architectures: Qwen2.5-VL [5] (3B, 7B, 32B, 72B), transformer-based models optimized for visual recognition; LLaVA-Next [14] (7B, 13B), integrating CLIP encoders with large language models; InternVL3 [27] (2B, 8B, 14B), using a ”ViT-MLP-LLM” architecture; the Mistral family , including Mistral-Small (24B) [3] and Pixtral variants (12B, 124B) [2]; and other architectures represented by Idefics2 (8B) [12] and Phi-3.5-vision (4B) [1]. All models were evaluated using publicly available checkpoints without task-specific fine-tuning to assess zero-shot capabilities.
@@
3.2 Evaluation Protocol
We evaluate model responses using a two-stage protocol:
Q1 (Binary Classification).
The model is asked whether the situation in the image is dangerous or safe. It selects one of two choices (A. Dangerous / B. Safe) and provides a brief reasoning. We compute precision, recall, and F1 score based on human-annotated binary ground truth labels.
Q2 (Open-ended Response).
For images classified as dangerous by the model and correctly aligned with the ground truth, we further ask what an appropriate emergency response should be. We evaluated these responses using GPT-4o as a judge, providing it with the image caption and our curated gold-standard answers as references. The judge was instructed to score each response on a scale from 0 to 1 based on its alignment with the reference materials and appropriateness for the emergency situation.
4 Results and Analysis
@@ 4.1 The Overreaction Phenomenon
To evaluate VLMs’ emergency recognition capabilities, we assessed their performance on both risk identification (Q1) and emergency response (Q2) tasks (Table 2). Our evaluation reveals a consistent pattern: models achieve high recall (0.70-1.00) in identifying dangerous situations, but precision is notably lower (0.51-0.72), indicating a systematic ”better-safe-than-sorry” bias—an ”overreaction problem.” This pattern persists across model families and parameter scales (2B-124B). Model size does not consistently correlate with improved precision; InternVL3 (8B) achieves the highest precision (0.72), outperforming larger variants. False positive rates (safe images misclassified as dangerous) ranged from 31% to 96%—substantially higher than false negative rates (missed emergencies), which ranged from 0% to 30%. Figure 3(a) illustrates this bias through the precision-recall tradeoff, with models clustering in the high-recall but lower-precision region.
Particularly concerning is that 10% of safe scenarios were misclassified by all 14 evaluated models, typically containing visual elements strongly associated with danger despite clear contextual safety cues. As shown in Figure 3(b), false positives consistently outnumber false negatives across all models. Increasing model size did not reliably mitigate this problem—within the Qwen2.5-VL family, precision improves modestly from 3B (0.51) to 72B (0.65), but recall fluctuates non-linearly. This persistent pattern across architectures and scales suggests the overreaction problem may be embedded in foundational visual understanding rather than reasoning capacity.
@@ 4.2 Emergency Response Evaluation
Beyond identifying emergencies, we evaluated how effectively models suggest appropriate actions for emergency situations. Models demonstrated moderate effectiveness (scores 0.46-0.70) in generating appropriate emergency responses for correctly identified dangerous scenarios.
Unlike risk identification, emergency response quality shows a clearer correlation with model size. We observe consistent scaling benefits: Qwen2.5-VL (0.46→0.70), InternVL3 (0.50→0.64), and Mistral (0.59→0.68). This linear relationship presents an interesting contrast to the overreaction problem: while contextual reasoning for risk assessment doesn’t consistently improve with scaling, procedural knowledge for emergency responses does. This suggests that protocols may be more amenable to parameter scaling than the nuanced contextual judgments required for accurate risk identification.
We found a disconnect between risk identification and response capabilities. InternVL3 (8B) achieved the highest precision in risk identification but only moderate response quality, while models with lower precision often produced better emergency responses. This suggests VLMs develop these capabilities through different mechanisms—procedural knowledge scaling more predictably than contextual reasoning. Even when models correctly identify emergencies, their ability to recommend appropriate interventions remains limited. Examples can be found in the supplementary material.
4.3 Category-Specific Analysis
Our analysis reveals measurable performance variations across emergency categories (Table 4). In risk identification (Q1), models generally performed best on Natural Disasters (ND, avg F1=0.75) compared to Accidents & Behaviors (AB, 0.73) and Personal Medical Emergencies (PME, 0.68). The PME category shows the highest variance, from exceptional (0.79 for Pixtral-Large) to extremely poor (0.41 for Phi-3.5-vision).
Model scaling effects vary by category. For PME, larger models generally show improved performance (e.g., Qwen2.5-VL improves from 0.68 to 0.73, Pixtral from 0.67 to 0.79), suggesting increased parameters help with subtle physiological cues. For AB and ND categories, we observe a non-linear relationship with size, where mid-sized models sometimes outperform larger counterparts (e.g., InternVL3-8B achieves 0.83 on AB, outperforming the 14B variant). The overreaction problem is most pronounced in the PME category (38% false positive rate) where models frequently misclassified training scenarios as genuine emergencies. Natural Disasters show the lowest false positive rate (27%), likely because environmental containment features provide more distinctive safety indicators.
These findings suggest VLMs struggle most with scenarios where emergency status depends on fine-grained contextual details rather than obvious visual elements. For emergency response quality, Q1 showed a clearer category preference (ND $>$ AB $>$ PME with differences up to 0.07), while Q2 exhibited smaller average differences between categories (approximately 0.03). This suggests that once a model correctly identifies an emergency, its ability to generate appropriate responses is more consistent across different emergency types. For detailed model-specific performance breakdowns across emergency categories, see the supplementary material.

### v3

@@ Abstract
Vision-Language Models (VLMs) have shown capabilities in interpreting visual content, but their reliability in safety-critical scenarios remains insufficiently explored. We introduce VERI, a diagnostic benchmark comprising 200 synthetic images (100 contrastive pairs) and additional 50 real-world images (25 pairs) for validation. Each emergency scene is paired with a visually similar but safe counterpart through human verification. Using a two-stage evaluation protocol (risk identification and emergency response), we assess 17 VLMs across medical emergencies, accidents, and natural disasters. Our analysis reveals an “overreaction problem”: models achieve high recall (70–100%) but suffer from low precision, misclassifying 31–96% of safe situations as dangerous. Seven safe scenarios were universally misclassified by all models. This “better-safe-than-sorry” bias stems from contextual overinterpretation (88–98% of errors). Both synthetic and real-world datasets confirm these systematic patterns, challenging VLM reliability in safety-critical applications. Addressing this requires enhanced contextual reasoning in ambiguous visual situations.
@@ 1 Introduction
Vision-Language Models (VLMs) have advanced from simple object recognition to sophisticated contextual scene understanding [18, 12, 4]. These capabilities now support diverse applications including content moderation, assistive technologies [28], and increasingly, safety-critical systems [19, 26]. However, their reliability in emergency scenarios remains insufficiently explored. A crucial question arises: can current VLMs reliably distinguish genuine emergencies from visually similar but safe situations? To investigate this systematically, we employ synthetic data enabling controlled comparisons while preserving visual similarity.
False visual perceptions of emergencies lead to costly resource mobilization and system failures. In October 2023, firefighters responded to incidents in New York and Seoul where high-definition fireplace videos were mistaken for actual fires [17, 24]. Beyond immediate costs (\$1,000-2,500 per incident), such false alarms create alarm fatigue, reducing trust in automated systems [20]. With annual false alarm costs exceeding \$1 billion in the United States alone and VLMs increasingly powering safety applications from smart home monitoring to CCTV surveillance [25], understanding their overreaction tendency becomes critical.
While prior work focused on domain-specific safety (medical diagnosis [8], autonomous driving [9], industrial robotics [10]), these approaches rely on specialized features and extensive domain training. We address a distinct challenge: everyday emergency recognition requiring general contextual reasoning. This demands distinguishing visually similar scenarios through subtle contextual cues rather than domain-specific patterns. Existing benchmarks evaluate models on isolated images without paired contrasts. They rarely test the critical ability to distinguish between visually similar scenarios where only context determines safety.
To address these gaps, we introduce VERI, a diagnostic benchmark of 200 synthetic images (100 contrastive pairs) created through multiple rounds of human verification to ensure controlled visual similarity and semantic distinction [^1] The VERI dataset is publicly available at
https://huggingface.co/datasets/Dasool/VERI-Emergency
. We validate our findings with 25 real-world contrastive pairs, confirming that the observed patterns generalize beyond synthetic data.
@@
  • We propose VERI with a two-stage evaluation protocol (risk identification and emergency response) to test VLMs in safety-critical scenarios.
  • We evaluate 17 VLMs (open-source and commercial), revealing a systematic “overreaction problem” where safe scenarios are frequently misclassified as dangerous, including 7 universally misclassified cases.
  • We analyze two recurring error patterns—contextual overinterpretation and visual misinterpretation—with contextual errors dominating (88–98
  • We show that the overreaction problem persists regardless of model scale, highlighting the need for targeted improvements in contextual emergency reasoning.
@@ 2.1 Dataset Design and Taxonomy
Effective emergency detection requires distinguishing between genuine threats and visually similar but safe situations. To address this challenge, the VERI dataset is designed to capture everyday risks rather than specialized domains such as medical diagnosis or industrial safety. The core design principle of VERI is the contrastive pair approach, where each entry consists of two images with high visual similarity but fundamentally different semantic implications: one depicting a genuine emergency requiring intervention and the other showing a visually similar but safe scenario. As illustrated in Figure 2, these pairs include contrasts such as actual medical emergencies versus training simulations, genuine accidents versus staged scenarios, and real disasters versus controlled environments. This structure allows models to be tested on their ability to distinguish superficially similar scenes with different safety implications. We organized VERI into three categories of emergency situations (Accidents & Unsafe Behaviors, Personal Medical Emergencies, and Natural Disasters), as shown in Table 1. Each category poses distinct visual challenges, ranging from subtle physiological cues to complex environmental contexts (Figure 2).
@@ 2.2 Image Creation Process
The use of synthetic images was necessitated by practical and ethical constraints: collecting real-world emergency images raises privacy and consent issues, while achieving the precise visual control required for contrastive pairs is virtually impossible with naturally occurring images. Moreover, synthetic data allows us to create perfectly matched pairs that isolate the specific contextual cues we aim to study, enabling more rigorous evaluation of VLMs’ contextual reasoning capabilities. We employed a multi-stage approach:
@@
Stage 3: Collaborative Refinement and Validation. To ensure diagnostic precision required for contrastive evaluation, we employed an iterative refinement approach. Three evaluators (two researchers and an independent annotator) assessed each image against predefined criteria: emergency scenes must clearly require intervention, safe scenes must be unambiguously non-threatening, and pairs must maintain visual similarity while being semantically distinct. When disagreements arose, images were modified using inpainting to address specific concerns, rather than accepting compromised examples through majority voting. This process continued until unanimous agreement was reached on all validation criteria across all evaluators ($n=200$ pairs, $600$ decisions), averaging 5+ refinement cycles per pair.
@@
Our process prioritized diagnostic value over quantity, focusing on expert-guided refinement to effectively challenge VLMs’ contextual reasoning. Each image pair required extensive manual refinement and creative input to achieve the visual precision necessary for meaningful contrastive evaluation. As the first benchmark for everyday emergency recognition, VERI emphasizes carefully curated diagnostic scenarios. This approach aligns with other influential diagnostic benchmarks in the literature, where controlled evaluation is more important than scale. For instance, the PAIRS benchmark comprises 200 images, and the study ’Judging LLM-as-a-Judge with MT-Bench and Chatbot Arena’ (Zheng et al., 2023) evaluated models using only 80 carefully constructed questions across eight fields.
@@ 2.3 Query Development
Commercial Models
Gemini-2.5-Flash $0.640$ $0.970$ $0.771$ $0.771$ 96
GPT-4o $0.645$ $0.980$ $0.778$ $0.670$ 98
Claude-4-Sonnet $0.580$ $0.910$ $0.708$ $0.737$ 91
@@ 3.1 Models
We evaluated 17 VLMs across different architectures, including both open-source models (2B-124B parameters) and commercial APIs. Open-source models included: Qwen2.5-VL [7] (3B, 7B, 32B, 72B), transformer-based models optimized for visual recognition; LLaVA-Next [16] (7B, 13B), integrating CLIP encoders with large language models; InternVL3 [29] (2B, 8B, 14B), using a ”ViT-MLP-LLM” architecture; the Mistral family, including Mistral-Small (24B) [3] and Pixtral variants (12B, 124B) [2]; and other architectures represented by Idefics2 (8B) [14] and Phi-3.5-vision (4B) [1]. Commercial models included GPT-4o [13], Gemini-2.5-Flash [12], and Claude-4-Sonnet [5].
3.2 Evaluation Protocol
We evaluate model responses using a two-stage protocol:
Q1 Binary Classification The model is asked whether the situation in the image is dangerous or safe. It selects one of two choices (A. Dangerous / B. Safe) and provides a brief reasoning. We compute precision, recall, and F1 score based on human-annotated binary ground truth labels.
Q2 Open-ended Response For images classified as dangerous by the model and correctly aligned with the ground truth, we further ask what an appropriate emergency response should be. We evaluated these responses using GPT-4o as a judge, providing it with the image caption and our curated gold-standard answers as references. The judge was instructed to score each response on a scale from 0 to 1 based on its alignment with the reference materials and appropriateness for the emergency situation. See Appendix B for detailed evaluation methodology.
@@ 3.3 The Overreaction Phenomenon
To evaluate VLMs’ emergency recognition capabilities, we assessed their performance on both risk identification (Q1) and emergency response (Q2) tasks (Table 2). Our evaluation reveals a consistent pattern: models achieve high recall (0.70–1.00) in identifying dangerous situations, but precision is notably lower (0.51–0.72), indicating a systematic “better-safe-than-sorry” bias, an “overreaction problem.” This pattern persists across both open-source and commercial models, with commercial APIs (GPT-4o: 0.645, Gemini-2.5-Flash: 0.640, Claude-4-Sonnet: 0.580) showing similar precision ranges to open-source counterparts. Model size does not consistently correlate with improved precision; InternVL3 (8B) achieves the highest precision among open-source models (0.721), while GPT-4o shows the best overall precision (0.645). False positive rates (safe images misclassified as dangerous) ranged from 31% to 96%, substantially higher than false negative rates (missed emergencies), which ranged from 2% to 30%. Figure 3(a) illustrates this bias through the precision–recall tradeoff, with models clustering in the high-recall but lower-precision region.
Particularly concerning is that 7 safe scenarios were misclassified by all 17 evaluated models, typically containing visual elements strongly associated with danger despite clear contextual safety cues. As shown in Figure 3(b), false positives consistently outnumber false negatives across all models. Even commercial models with sophisticated training exhibit this bias, with GPT-4o achieving a 38% false positive rate, Gemini-2.5-Flash 39%, and Claude-4-Sonnet 42%. This persistent pattern across architectures, scales, and development approaches suggests that the overreaction problem may be embedded in foundational visual understanding rather than in higher-level reasoning capacity.
@@ 3.4 Emergency Response Evaluation
Beyond identifying emergencies, we evaluated how effectively models suggest appropriate actions for emergency situations. Models demonstrated moderate effectiveness (scores 0.46–0.77) in generating appropriate emergency responses for correctly identified dangerous scenarios.
Unlike risk identification, emergency response quality shows clearer correlation with model size within families. We observe consistent scaling benefits: Qwen2.5-VL (0.46→0.70), InternVL3 (0.50→0.64), and Mistral (0.59→0.68). Commercial models achieved higher scores (Gemini-2.5-Flash: 0.771, Claude-4-Sonnet: 0.737, GPT-4o: 0.670). This scaling relationship contrasts with the overreaction problem: while contextual reasoning for risk assessment does not consistently improve with scaling, procedural knowledge for emergency responses does.
We found a disconnect between risk identification and response capabilities. InternVL3 (8B) achieved the highest precision in risk identification among open-source models but only moderate response quality (0.610), while models with lower precision often produced better emergency responses. This suggests VLMs develop these capabilities through different mechanisms, with procedural knowledge scaling more predictably than contextual reasoning.
To validate GPT-4o evaluation reliability, we conducted inter-judge analysis with Claude-4.1-Sonnet (N=300) [6], showing substantial agreement ($r=0.719$, $p<0.001$), confirming the robustness of our Q2 scores.
@@
Commercial Models
Gemini-2.5-Flash 0.805 0.759 0.753
GPT-4o 0.779 0.778 0.777
Claude-4-Sonnet 0.686 0.708 0.725
@@
3.5 Category-Specific Analysis
Our analysis reveals performance variations across emergency categories (Table 4). Models generally performed best on Natural Disasters (ND, avg F1=0.75) compared to Accidents & Behaviors (AB, 0.73) and Personal Medical Emergencies (PME, 0.69), with PME showing the highest variance (0.408-0.805). Model scaling effects vary by category: in PME, larger models generally improve (Qwen2.5-VL: 0.681→0.727, Pixtral: 0.667→0.790), possibly due to better recognition of subtle physiological cues, while in AB and ND, mid-sized models sometimes outperform larger variants—InternVL3-8B achieved 0.825 on AB versus 0.761 for the 14B model. Commercial models showed robust performance across all categories (Gemini-2.5-Flash: 0.805 PME, GPT-4o: 0.777-0.779).
3.6 Real-World Validation
To validate whether our findings generalize beyond synthetic data, we collected 25 real-world contrastive pairs (50 images) across the same three categories. These images were evaluated using the same Q1 protocol to test whether the overreaction pattern persists in natural images. Finding closely matched real-world pairs was challenging given our controlled emergency-safe scenarios.
Table 5 shows that results confirm the same pattern observed on the synthetic benchmark: models achieve high recall (88-100%) but lower precision (57-68%), with false positives consistently dominating. The consistent overreaction bias across both synthetic and real-world images validates the diagnostic value of our synthetic approach. See Appendix A for detailed examples and analysis.
Synthetic (VERI) Real-World (25 Pairs)
Model P R F1 P R F1
Claude-4-Sonnet 0.580 0.910 0.708 0.568 1.000 0.725
GPT-4o 0.645 0.980 0.778 0.677 0.920 0.780
Qwen2.5-VL-72B 0.652 0.900 0.756 0.647 0.880 0.746
Pixtral-Large 0.632 0.980 0.769 0.649 0.960 0.774
Table 5: Synthetic vs. real-world performance comparison showing consistent high recall but lower precision.
3.7 Prompt Sensitivity Analysis
To test whether simple instruction tweaks can mitigate overreaction, we re-ran Q1 with a conservative prompt that explicitly says: “Choose Dangerous only if there is a clear and immediate risk requiring action; otherwise answer Safe.” and forces the output to A/B. Figure 5 summarizes the effect across four representative models (Qwen2.5-VL-72B, Pixtral-Large, InternVL3-14B, Claude-4.0). Across the board, precision rises while recall drops, i.e., false positives decrease but false negatives increase. This confirms a clear trade-off: conservative phrasing can make models less trigger-happy, but at the cost of missing more true emergencies. Hence, prompt tuning alone cannot resolve the overreaction problem without sacrificing coverage. Similarly, post-hoc threshold optimization shows the same limitation—even at $F_{1}$-optimal thresholds, false positives dominate (Appendix C), indicating that architectural interventions are needed beyond simple calibration.
Figure 5: Prompt sensitivity on Q1. For each model, arrows point from the default prompt (open marker) to a conservative prompt (filled marker). Precision consistently increases (top), whereas recall decreases (bottom), demonstrating a precision–recall trade-off induced by conservative phrasing.